OpenStudy (anonymous):

if f(x)=squareroot of x+3 and g(x)=2xsqaured then find g0f (-1)

OpenStudy (anonymous):

gf(x) = 2 * [sqrt(x+3)]^2 = 2*(x+3) gf(-1) = 2 *(-1+3) = 4

OpenStudy (anonymous):

I am writing it down to see if I can understand it

OpenStudy (anonymous):

Thank you so very much!! I greatly appreciate the help!! GOD bless you for your kindness<><